OpenStudy (anonymous):

Evaluate ( 5.0 x 10^17 )( 8.0 x 10^-17 ). Be sure to follow the guidelines for entering answers in scientific notation.

OpenStudy (anonymous):

yupyup so what do you think? clue -17+17 != 1 :D

OpenStudy (anonymous):

@RONNCC is that evaluated in scientific notation

OpenStudy (anonymous):

yesh that is part of this problem. clue. you can just multiply. why? because you're just doing 4 multiplications. :P 5e17 = (5)(10^17)

OpenStudy (anonymous):

oh okay so is 1 my answer to the problem

OpenStudy (zehanz):

What about the 5x8?

OpenStudy (anonymous):

@ZeHanz 5x8 =40

OpenStudy (zehanz):

So you now have 40 x 10^0. That is almost scientific notation. In s.n. the first number (here 40) should always be between 1 and 10...

OpenStudy (anonymous):

so that means 400 is the answer

OpenStudy (zehanz):

No, it was 40 x 10^0. Now 10^0 = 1 , so the number is 40 x 1 =40. Still, you are not done, because you have to use scientific notation for the number 40. Because scientific notation is always: a x 10^b, where a is between 1 and 10. 40 = 4 x 10, so to make this scientific notation, just write 10 as a power of 10: 10^1. All in all, the outcome is 4 x 10^1, or, probably better, because of the decimals in the original numbers: 4.0 x 10^1. And no, this is not a handy way to write down the number 40, it is just meant as an exercise... ;)